# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ Appendix B Parton construction
Here $f_{r}$ and $\Phi_{r}$ are fermionic and bosonic partons. We take the boson to have the physical electric charge and the fermions to be electrically neutral. As usual these partons are coupled to a dynamical $\mathrm{U}(1)$ gauge field. Both $f$ and $\Phi$ will be at the same mean filling $\nu$ as the electrons. Now consider a state where the bosons have formed a gapped $\mathbb{Z}_{q}$ topologically ordered state. A simple option for the fermions is that they form a Fermi surface. The resulting state is a gapless incompressible insulator with a neutral Fermi surface which is closely analogous to the spinon Fermi surface of spinful electronic Mott insulator at half-filling. It is readily seen that the low energy theory of this state is described by a field theory with the structure
@@
We assign physical charge $1/3$ to each of these 3 fermions. This representation comes with an $SU(3)$ gauge redundancy. Now consider a ‘Higgs’ condensate where
$T_{ab}=\langle d^{\dagger}_{ra}d_{rb}\rangle~{}~{}\neq 0$ (18)
such that the SU(3) is broken down to $\mathbb{Z}_{3}$. Then the low energy theory will be a deconfined $\mathbb{Z}_{3}$ gauge theory where the elementary gauge charge is a fermion, and the gauge flux is a boson. The Higgs condensation also implies that the three fermion species $d_{a}$ (with $a=1,2,3$) can all be identified with each other, and therefore there is just a single fermion species. These $d$-fermions are then at integer filling and can form a band insulator. Thus we get a gapped deconfined $\mathbb{Z}_{3}$ gauge theory with fermionic charges that carry physical electric charge $1/3$.
@@ Appendix C Results about Abelian TOs
In this section we will prove a number of results about Abelian TOs. Most of the proofs rely only on basic facts from group theory, i.e. the first isomorphism theorem and Lagrange’s theorem. They are as self-contained as possible. We will later find these proofs useful even when we consider non-Abelian TOs, $\mathcal{C}$, as we will be able to make arguments to apply them to the subset of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$. Moreover, once we have proven that all minimal TOs are Abelian all of these results will be able to be used in full to constrain the possible Abelian TOs.
The structure of this appendix is as follows. In App. C.1 we introduce our group theoretic approach by discussing condensation in an Abelian TO. In App. C.2 we introduce a single symmetry, time reversal, and give some results about the size of a time reversal invariant Abelian TO. In App. C.3 we introduce a second symmetry, and consider Abelian SETs of $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$.
@@ C.1 Results about Abelian TO and anyon condensation
Let $\mathcal{A}$ be a (bosonic or fermionic) parent Abelian TO. We suppose that there is a boson $v\in\mathcal{A}$ which has order $n_{v}$. We then want to form a child TO $\mathcal{A}^{\prime}$ from $\mathcal{A}$ by condensing $v$. We will give a more precise definition of condensation and show that $|\mathcal{A}|=n_{v}^{2}|\mathcal{A}^{\prime}|$. Much of this subsection follows ideas discussed in Burnell (2018), but we reformulate them in terms of group theoretic language to build fluency. The tools developed here will prove useful once we introduce symmetries.
@@
The braiding phase $\theta\colon\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{S}^{1}$ is a symmetric bilinear form, i.e. it has the properties that
@@
for all $a,b,d\in\mathcal{A}$.
@@
We call $\mathcal{A}$ a bosonic TO if braiding is non-degenerate on $\mathcal{A}$, i.e. if $\theta_{a,b}=0$ for all $b\in\mathcal{A}$ implies that $a=1$, where $1$ is a boson. We call $\mathcal{A}$ a fermionic TO if $\theta_{a,b}=0$ for all $b\in\mathcal{A}$ implies that $a=1,c$ where $c^{2}=1$ and $\theta_{c}=\pi$, i.e. if braiding is non-degenerate up to the fundamental fermion.
@@
Let $\mathcal{A}$ be an Abelian TO. Then take $\mathcal{S}\subseteq\mathcal{A}$ to be a subgroup of $\mathcal{A}$. We define the complement of $\mathcal{S}$ to be given by:
@@
We will use a result from character theory to establish this. First we recall that a character $\chi$ of a finite Abelian group $\mathcal{A}$ is a homomorphism $\chi\colon\mathcal{A}\rightarrow\mathbb{S}^{1}$. The set of characters of $\mathcal{A}$ is labeled as $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, where this can be treated as an Abelian group under pointwise addition, i.e. $(\chi_{1}+\chi_{2})(a)=\chi_{1}(a)+\chi_{2}(a)$ for $\chi_{1},\chi_{2}\in\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ and $a\in\mathcal{A}$. A result from character theory tells us that the number of distinct characters of a finite Abelian group $\mathcal{A}$ is precisely $|\mathcal{A}|$ Conrad .
We know that $\theta_{a,\cdot}\colon\mathcal{A}\rightarrow\mathbb{S}^{1}$ is a homomorphism for every $a\in\mathcal{A}$. Further, because $\mathcal{A}$ is a bosonic TO the braiding phase is non-degenerate. Thus if $a,b\in\mathcal{A}$ and $a\neq b$ then there exists some $d\in\mathcal{A}$ such that $\theta_{a,d}\neq\theta_{b,d}$. Then we see that $\theta_{a,\cdot}=\theta_{b,\cdot}$ as an element of $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ if, and only if, $a=b$. This means that we have found $|\mathcal{A}|$ distinct elements of $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, and by the logic above we see that for every character $\chi\in\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ there exists one, and only one, $a\in\mathcal{A}$ such that $\chi=\theta_{a,\cdot}$.
We now want to consider a function $B\colon\mathcal{A}\rightarrow\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$, i.e. $B(a)\colon\mathcal{S}\rightarrow\mathbb{S}^{1}$ is a homomorphism for every $a\in\mathcal{A}$. We define it as $B(a)(b)=\theta_{a,b}$, where $b$ is restricted to lie in the subgroup $\mathcal{S}$. We note that $B(ad)=B(a)+B(d)$ for $a,d\in\mathcal{A}$ so $B$ is a homomorphism from $\mathcal{A}$ to $\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$. Now suppose that $a\in\mathrm{Ker}(B)$, then $B(a)(b)=0$ for all $b\in\mathcal{S}$. But then we see that $\mathrm{Ker}(B)\simeq\mathcal{S}^{\perp}$. Now consider $\chi\in\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})\subseteq\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, then by the logic above we know that there is some $a\in\mathcal{A}$ such that $\chi=B(a)$. Thus $\mathrm{Im}(B)\simeq\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$. The first isomorphism theorem tells us that
@@
$|\mathcal{A}|=|\mathcal{S}||\mathcal{S}^{\perp}|\cdot\begin{cases}1&\mbox{if }c\notin\mathcal{S}\\ 1/2&\mbox{if }c\in\mathcal{S}\end{cases}.$ (23)
@@
Let $\mathcal{A}$ be a bosonic Abelian TO with two subgroups $\mathcal{S}$ and $\mathcal{R}$. Further suppose that braiding is non-degenerate when restricted to these subgroups and that $\theta_{r,s}=0$ for all $r\in\mathcal{R}$ and all $s\in\mathcal{S}$. Then we say that $\mathcal{A}$ factorizes as a TO and write $\mathcal{A}=\mathcal{S}\boxtimes\mathcal{R}$.
@@
Suppose $\mathcal{A}$ is a bosonic Abelian TO, and $\mathcal{S}$ is a subgroup. If the braiding restricted to $\mathcal{S}$ is non-degenerate, then $\mathcal{A}=\mathcal{S}\boxtimes\mathcal{S}^{\perp}$.
@@
First, we show that the braiding on $\mathcal{S}^{\perp}$ is also non-degenerate. If it is not, then there exists $x\in\mathcal{S}^{\perp},x\neq 1$ such that $x\in(\mathcal{S}^{\perp})^{\perp}$. However, by Collorary C.9 $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$, thus $x\in\mathcal{S}$ as well, which contradicts the assumption that $\mathcal{S}$ has non-degenerate braiding. The argument also shows that $\mathcal{S}\cap\mathcal{S}^{\perp}=\{1\}$.
@@
The theorem is a special case of a more general result about factorization of modular tensor categories (MTC), Theorem 3.13 in Drinfeld et al. (2010). Namely, if a MTC $\mathcal{C}$ has a modular subcategory $\mathcal{B}$, then the MTC can be factorized as $\mathcal{C}=\mathcal{B}\boxtimes\mathcal{B}^{\perp}$, where $\mathcal{B}^{\perp}$ is the “complement” of $\mathcal{B}$ in $\mathcal{C}$.
@@
First suppose that $\mathcal{A}$ is a bosonic Abelian TO. Since $v$ is a boson it braids trivially with itself, and thus will with all of its powers. So we have that $\langle v\rangle$ is a subgroup of $\langle v\rangle^{\perp}\subseteq\mathcal{A}$. Thus $\mathcal{A}^{\prime}=\langle v\rangle^{\perp}/\langle v\rangle$ is well defined. Now we use Cor. C.9 to see that $(\langle v\rangle^{\perp})^{\perp}=\langle v\rangle$. Thus if $a\in\mathcal{A}$ braids trivially with every element of $\langle v\rangle^{\perp}$ it must lie in $\langle v\rangle$. Since in $\mathcal{A}^{\prime}$ every element of $\langle v\rangle$ is identified with the vacuum, we must then have that braiding is non-degenerate on $\mathcal{A}^{\prime}$. Thus $\mathcal{A}^{\prime}$ is a bosonic TO.
@@
Suppose now that $\mathcal{A}$ is a fermionic Abelian TO. We first note that since $v$ is a boson no power of $v$ can be a fermion and hence $c\notin\langle v\rangle$. For the same reasons as above we have that $\mathcal{A}^{\prime}=\langle v\rangle^{\perp}/\langle v\rangle$ is well-defined. By Cor. C.9 we see that $(\langle v\rangle^{\perp})^{\perp}=\langle v\rangle\times\{1,c\}$. Then we see that braiding is non-degenerate on $\mathcal{A}^{\prime}$ up to the fundamental fermion $c$. Thus $\mathcal{A}^{\prime}$ is a fermionic TO. By Lem. 23 we see that $|\mathcal{A}|=|\langle v\rangle||\langle v\rangle^{\perp}|=n_{v}^{2}|\mathcal{A}^{\prime}|$. ∎
@@
A simple example of this is given by $\mathbb{Z}_{n}$ gauge theory. There we have a boson $e\in\mathcal{A}$ with order $n$. If we condense $e$ then the resulting theory is trivial, i.e. $\mathcal{A}^{\prime}=\{1\}$. We then have that $|\mathcal{A}|=n^{2}|\mathcal{A}^{\prime}|$.
@@ C.2 Time reversal invariant Abelian TOs
A particularly powerful algebraic result about time reversal invariant Abelian bosonic TOs was given in Lee and Tachikawa (2018). We state this result as a theorem, though we must first give a definition of time-reversal invariant Abelian TOs. At the end of this subsection we extend the results of Lee and Tachikawa (2018) to include fermionic Abelian TOs as well as strengthen the bounds on anyon count for Abelian bosonic TOs.
@@
Let $\mathcal{A}$ be an Abelian TO. We call $\mathcal{A}$ a time-reversal invariant Abelian TO if it has an automorphism $\mathcal{T}\colon\mathcal{A}\rightarrow\mathcal{A}$ with the following properties. First, $\mathcal{T}^{2}$ is the identity on $\mathcal{A}$. Also, if $a\in\mathcal{A}$ then $\theta_{\mathcal{T}a}=-\theta_{a}$ so time reversal takes the topological spin to its negative. Finally, if $\mathcal{A}$ is a fermionic TO then $\mathcal{T}c=c$, i.e. time reversal sends the fundamental fermion to itself.
@@
In this section we have not specified whether or not $c$ is a Kramers doublet under $\mathcal{T}$, though the anomaly structure of our theory will depend on this Lapa and Levin (2019). We further note that other studies of the action of time reversal on Abelian TOs have allowed $\mathcal{T}^{2}=C$, where $C$ is charge conjugation Delmastro and Gomis (2021). However, this action of time reversal is not consistent with $\mathrm{U}(1)$ charge symmetry. Thus we will restrict $\mathcal{T}^{2}$ to be the identity on $\mathcal{A}$.
@@
This proof follows immediately from (3.11) of Lee and Tachikawa (2018), which showed that $|\mathcal{A}_{b}|=l^{2}$ for some integer $l$.
We will review the derivation of $|\mathcal{A}_{b}|=l^{2}$ for clarity, since only part of the discussion in Lee and Tachikawa (2018) is necessary for our purposes. Define:
@@
where $\overline{a}$ is the dual of $a$. Suppose that $a\in\mathrm{Im}(f)$, i.e. $a=b\times\mathcal{T}b$ for some $b\in\mathcal{A}_{b}$. Then we see that $\mathcal{T}a=a$ since $\mathcal{T}$ squares to the identity. Thus we see that $a\in\mathrm{Ker}(g)$ and hence that $\mathrm{Im}(f)$ is a subgroup of $\mathrm{Ker}(g)$. We can thus define:
@@
We will now use the fact that $\theta_{\mathcal{T}a,\mathcal{T}b}=-\theta_{a,b}$ for all $a,b\in\mathcal{A}_{b}$. Note that this means that $\theta_{a,\mathcal{T}b}=-\theta_{\mathcal{T}a,b}=\theta_{\mathcal{T}\overline{a},b}$. If we add $\theta_{a,b}$ to both sides then we can use its bilinearity to see that $\theta_{a,b\times\mathcal{T}b}=\theta_{a\times\mathcal{T}\overline{a},b}$, i.e. that $\theta_{a,f(b)}=\theta_{g(a),b}$. Then consider $a\in\mathrm{Ker}(g)$ we see that $0=\theta_{1,b}=\theta_{g(a),b}=\theta_{a,f(b)}$ for every $b\in\mathcal{A}_{b}$. Similarly we see that for $b\in\mathrm{Ker}(f)$ we have $0=\theta_{g(a),b}$ for all $a\in\mathcal{A}_{b}$. We thus have
@@
$|\mathrm{Ker}(g)|$ $\leq|\mathrm{Im}(f)|^{\perp}\text{ by above}$ (29)
@@
By Cor. C.9 we know that $(\mathrm{Im}(f)^{\perp})^{\perp}=\mathrm{Im}(f)$. Thus the only elements $a\in\mathcal{A}_{b}$ that braid trivially with all elements of $\mathrm{Im}(f)^{\perp}$ are those in $\mathrm{Im}(f)$. This means that since $\mathcal{C}=\mathrm{Im}(f)^{\perp}/\mathrm{Im}(f)$ we must have that braiding is non-degenerate on $\mathcal{C}$.

### v2

@@ Appendix B Parton construction
Here $f_{r}$ and $\Phi_{r}$ are fermionic and bosonic partons. We take the boson to have the physical electric charge and the fermions to be electrically neutral. As usual these partons are coupled to a dynamical $\mathrm{U}(1)$ gauge field. Both $f$ and $\Phi$ will be at the same mean filling $\nu$ as the electrons. Now consider a state where the bosons have formed a gapped $\mathbb{Z}_{q}$ topologically ordered state. A simple option for the fermions is that they form a Fermi surface. The resulting state is a gapless incompressible insulator with a neutral Fermi surface which is closely analogous to the spinon Fermi surface of spinful electronic Mott insulator at half filling. It is readily seen that the low-energy theory of this state is described by a field theory with the structure
@@
We assign physical charge $1/3$ to each of these 3 fermions. This representation comes with an $\mathrm{SU}(3)$ gauge redundancy. Now consider a ‘Higgs’ condensate where
$T_{ab}=\langle d^{\dagger}_{ar}d_{br}\rangle~{}~{}\neq 0$ (18)
such that the $\mathrm{SU}(3)$ is broken down to $\mathbb{Z}_{3}$. Then the low-energy theory will be a deconfined $\mathbb{Z}_{3}$ gauge theory where the elementary gauge charge is a fermion, and the gauge flux is a boson. The Higgs condensation also implies that the three fermion species $d_{a}$ (with $a=1,2,3$) can all be identified with each other, and therefore there is just a single fermion species. These $d$-fermions are then at integer filling and can form a band insulator. Thus we get a gapped deconfined $\mathbb{Z}_{3}$ gauge theory with fermionic charges that carry physical electric charge $1/3$.
@@ Appendix C Results about Abelian TOs
In this section we prove a number of results about Abelian TOs. Most of the proofs rely only on basic facts from group theory i.e., the first isomorphism theorem and Lagrange’s theorem. They are as self-contained as possible. We later find these proofs useful even when we consider non-Abelian TOs $\mathcal{C}$ because we will be able to make arguments to apply them to the subset of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$. Moreover, once we have proven that all minimal TOs are Abelian all of these results will be able to be used in full to constrain the possible Abelian TOs.
The structure of this appendix is as follows. In Appendix C.1 we introduce our group-theoretic approach by discussing condensation in an Abelian TO. In Appendix C.2 we introduce a single symmetry, time reversal, and give some results about the size of a time reversal invariant Abelian TO. In Appendix C.3 we introduce a second symmetry, and consider Abelian SETs of $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$.
@@ C.1 Results about Abelian TO and anyon condensation
Let $\mathcal{A}$ be a (bosonic or fermionic) parent Abelian TO. We suppose that there is a boson $v\in\mathcal{A}$ which has order $n_{v}$. We then want to form a child TO $\mathcal{A}^{\prime}$ from $\mathcal{A}$ by condensing $v$. We give a more precise definition of condensation and show that $|\mathcal{A}|=n_{v}^{2}|\mathcal{A}^{\prime}|$. Much of this section follows ideas discussed in Ref. Burnell (2018), but we reformulate them in terms of group-theoretic language to build fluency. The tools developed here will prove useful once we introduce symmetries.
@@
The braiding phase $\theta\colon\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{S}^{1}$ is a symmetric bilinear form i.e., it has the properties that
@@
for all $a,b,d\in\mathcal{A}$. Additionally $\theta_{a,a}=2\theta_{a}$, where $\theta_{a}$ is the topological spin of $a\in\mathcal{A}$.
@@
We call $\mathcal{A}$ a bosonic TO if braiding is nondegenerate on $\mathcal{A}$ i.e., if $\theta_{a,b}=0$ for all $b\in\mathcal{A}$ implies that $a=1$, where $1$ is a boson. We call $\mathcal{A}$ a fermionic TO if $\theta_{a,b}=0$ for all $b\in\mathcal{A}$ implies that $a=1,c$ where $c^{2}=1$ and $\theta_{c}=\pi$ i.e., if braiding is nondegenerate up to the fundamental fermion.
@@
Let $\mathcal{A}$ be an Abelian TO. Then take $\mathcal{S}\subseteq\mathcal{A}$ to be a subgroup of $\mathcal{A}$. We define the complement of $\mathcal{S}$ to be given by
@@
We use a result from character theory to establish this. First we recall that a character $\chi$ of a finite Abelian group $\mathcal{A}$ is a homomorphism $\chi\colon\mathcal{A}\rightarrow\mathbb{S}^{1}$. The set of characters of $\mathcal{A}$ is labeled as $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, where this can be treated as an Abelian group under pointwise addition i.e., $(\chi_{1}+\chi_{2})(a)=\chi_{1}(a)+\chi_{2}(a)$ for $\chi_{1},\chi_{2}\in\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ and $a\in\mathcal{A}$. A result from character theory tells us that the number of distinct characters of a finite Abelian group $\mathcal{A}$ is precisely $|\mathcal{A}|$ Conrad .
We know that $\theta_{a,\cdot}\colon\mathcal{A}\rightarrow\mathbb{S}^{1}$ is a homomorphism for every $a\in\mathcal{A}$. Furthermore, because $\mathcal{A}$ is a bosonic TO the braiding phase is nondegenerate. Thus if $a,b\in\mathcal{A}$ and $a\neq b$ then there exists some $d\in\mathcal{A}$ such that $\theta_{a,d}\neq\theta_{b,d}$. Then we see that $\theta_{a,\cdot}=\theta_{b,\cdot}$ as an element of $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ if, and only if, $a=b$. This means that we have found $|\mathcal{A}|$ distinct elements of $\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, and by the logic above we see that for every character $\chi\in\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$ there exists one, and only one, $a\in\mathcal{A}$ such that $\chi=\theta_{a,\cdot}$.
We now want to consider a function $B\colon\mathcal{A}\rightarrow\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$ i.e., $B(a)\colon\mathcal{S}\rightarrow\mathbb{S}^{1}$ is a homomorphism for every $a\in\mathcal{A}$. We define it as $B(a)(b)=\theta_{a,b}$, where $b$ is restricted to lie in the subgroup $\mathcal{S}$. We note that $B(ad)=B(a)+B(d)$ for $a,d\in\mathcal{A}$ so $B$ is a homomorphism from $\mathcal{A}$ to $\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$. Now suppose that $a\in\mathrm{Ker}(B)$, then $B(a)(b)=0$ for all $b\in\mathcal{S}$. But then we see that $\mathrm{Ker}(B)\simeq\mathcal{S}^{\perp}$. Now consider $\chi\in\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})\subseteq\mathrm{Hom}(\mathcal{A},\mathbb{S}^{1})$, then by the logic above we know that there is some $a\in\mathcal{A}$ such that $\chi=B(a)$. Thus $\mathrm{Im}(B)\simeq\mathrm{Hom}(\mathcal{S},\mathbb{S}^{1})$. The first isomorphism theorem tells us that
@@
$|\mathcal{A}|=|\mathcal{S}||\mathcal{S}^{\perp}|\times\begin{cases}1&\mbox{if }c\notin\mathcal{S}\\ 1/2&\mbox{if }c\in\mathcal{S}\end{cases}.$ (23)
@@
Let $\mathcal{A}$ be a bosonic Abelian TO with two subgroups $\mathcal{S}$ and $\mathcal{R}$. Further suppose that braiding is nondegenerate when restricted to these subgroups and that $\theta_{r,s}=0$ for all $r\in\mathcal{R}$ and all $s\in\mathcal{S}$. Then we say that $\mathcal{A}$ factorizes as a TO and write $\mathcal{A}=\mathcal{S}\boxtimes\mathcal{R}$.
@@
Suppose $\mathcal{A}$ is a bosonic Abelian TO, and $\mathcal{S}$ is a subgroup. If the braiding restricted to $\mathcal{S}$ is nondegenerate, then $\mathcal{A}=\mathcal{S}\boxtimes\mathcal{S}^{\perp}$.
@@
First, we show that the braiding on $\mathcal{S}^{\perp}$ is also nondegenerate. If it is not, then there exists $x\in\mathcal{S}^{\perp},x\neq 1$ such that $x\in(\mathcal{S}^{\perp})^{\perp}$. However, by Collorary C.9 $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$, thus $x\in\mathcal{S}$ as well, which contradicts the assumption that $\mathcal{S}$ has nondegenerate braiding. The argument also shows that $\mathcal{S}\cap\mathcal{S}^{\perp}=\{1\}$.
@@
The theorem is a special case of a more general result about factorization of modular tensor categories (MTC), Theorem 3.13 in Ref. Drinfeld et al. (2010). Namely, if a MTC $\mathcal{C}$ has a modular subcategory $\mathcal{B}$, then the MTC can be factorized as $\mathcal{C}=\mathcal{B}\boxtimes\mathcal{B}^{\perp}$, where $\mathcal{B}^{\perp}$ is the “complement” of $\mathcal{B}$ in $\mathcal{C}$.
@@
First suppose that $\mathcal{A}$ is a bosonic Abelian TO. Since $v$ is a boson it braids trivially with itself, and thus will with all of its powers. So we have that $\langle v\rangle$ is a subgroup of $\langle v\rangle^{\perp}\subseteq\mathcal{A}$. Thus $\mathcal{A}^{\prime}=\langle v\rangle^{\perp}/\langle v\rangle$ is well defined. Now we use Cor. C.9 to see that $(\langle v\rangle^{\perp})^{\perp}=\langle v\rangle$. Thus if $a\in\mathcal{A}$ braids trivially with every element of $\langle v\rangle^{\perp}$ it must lie in $\langle v\rangle$. Since in $\mathcal{A}^{\prime}$ every element of $\langle v\rangle$ is identified with the vacuum, we must then have that braiding is nondegenerate on $\mathcal{A}^{\prime}$. Thus $\mathcal{A}^{\prime}$ is a bosonic TO.
@@
Suppose now that $\mathcal{A}$ is a fermionic Abelian TO. We first note that since $v$ is a boson no power of $v$ can be a fermion and hence $c\notin\langle v\rangle$. For the same reasons as above we have that $\mathcal{A}^{\prime}=\langle v\rangle^{\perp}/\langle v\rangle$ is well-defined. By Cor. C.9 we see that $(\langle v\rangle^{\perp})^{\perp}=\langle v\rangle\times\{1,c\}$. Then we see that braiding is nondegenerate on $\mathcal{A}^{\prime}$ up to the fundamental fermion $c$. Thus $\mathcal{A}^{\prime}$ is a fermionic TO. By Lem. 23 we see that $|\mathcal{A}|=|\langle v\rangle||\langle v\rangle^{\perp}|=n_{v}^{2}|\mathcal{A}^{\prime}|$. ∎
@@
A simple example of this is given by $\mathbb{Z}_{n}$ gauge theory. There we have a boson $e\in\mathcal{A}$ with order $n$. If we condense $e$ then the resulting theory is trivial i.e., $\mathcal{A}^{\prime}=\{1\}$. We then have that $|\mathcal{A}|=n^{2}|\mathcal{A}^{\prime}|$.
@@ C.2 Time reversal invariant Abelian TOs
A particularly powerful algebraic result about time reversal invariant Abelian bosonic TOs was given in Ref. Lee and Tachikawa (2018). We state this result as a theorem, although we must first give a definition of time-reversal invariant Abelian TOs. At the end of this section we extend the results of Ref. Lee and Tachikawa (2018) to include fermionic Abelian TOs as well as strengthen the bounds on anyon count for Abelian bosonic TOs.
@@
Let $\mathcal{A}$ be an Abelian TO. We call $\mathcal{A}$ a time-reversal invariant Abelian TO if it has an automorphism $\mathcal{T}\colon\mathcal{A}\rightarrow\mathcal{A}$ with the following properties. First, $\mathcal{T}^{2}$ is the identity on $\mathcal{A}$. Also, if $a\in\mathcal{A}$ then $\theta_{\mathcal{T}a}=-\theta_{a}$ so time reversal takes the topological spin to its negative. Finally, if $\mathcal{A}$ is a fermionic TO then $\mathcal{T}c=c$ i.e., time reversal sends the fundamental fermion to itself.
@@
In this section we have not specified whether $c$ is a Kramers doublet under $\mathcal{T}$, although the anomaly structure of our theory will depend on this Lapa and Levin (2019). We further note that other studies of the action of time reversal on Abelian TOs have allowed $\mathcal{T}^{2}=C$, where $C$ is charge conjugation Delmastro and Gomis (2021). However, this action of time reversal is not consistent with $\mathrm{U}(1)$ charge symmetry. Thus we restrict $\mathcal{T}^{2}$ to be the identity on $\mathcal{A}$.
@@
This proof follows immediately from (3.11) of Ref. Lee and Tachikawa (2018), which showed that $|\mathcal{A}_{b}|=l^{2}$ for some integer $l$.
We will review the derivation of $|\mathcal{A}_{b}|=l^{2}$ for clarity, since only part of the discussion in Ref. Lee and Tachikawa (2018) is necessary for our purposes. Define:
@@
where $\overline{a}=a^{-1}$ is the dual of $a$. Suppose that $a\in\mathrm{Im}(f)$ i.e., $a=b\times\mathcal{T}b$ for some $b\in\mathcal{A}_{b}$. Then we see that $\mathcal{T}a=a$ since $\mathcal{T}$ squares to the identity. Thus we see that $a\in\mathrm{Ker}(g)$ and hence that $\mathrm{Im}(f)$ is a subgroup of $\mathrm{Ker}(g)$. We can thus define:
@@
We now use the fact that $\theta_{\mathcal{T}a,\mathcal{T}b}=-\theta_{a,b}$ for all $a,b\in\mathcal{A}_{b}$. Note that this means that $\theta_{a,\mathcal{T}b}=-\theta_{\mathcal{T}a,b}=\theta_{\mathcal{T}\overline{a},b}$. If we add $\theta_{a,b}$ to both sides then we can use its bilinearity to see that $\theta_{a,b\times\mathcal{T}b}=\theta_{a\times\mathcal{T}\overline{a},b}$ i.e., that $\theta_{a,f(b)}=\theta_{g(a),b}$. Then consider $a\in\mathrm{Ker}(g)$ we see that $0=\theta_{1,b}=\theta_{g(a),b}=\theta_{a,f(b)}$ for every $b\in\mathcal{A}_{b}$. Similarly we see that for $b\in\mathrm{Ker}(f)$ we have $0=\theta_{g(a),b}$ for all $a\in\mathcal{A}_{b}$. We thus have
@@
$|\mathrm{Ker}(g)|$ $\leq|\mathrm{Im}(f)^{\perp}|\text{ by above}$ (29)
@@
By Cor. C.9 we know that $(\mathrm{Im}(f)^{\perp})^{\perp}=\mathrm{Im}(f)$. Thus the only elements $a\in\mathcal{A}_{b}$ that braid trivially with all elements of $\mathrm{Im}(f)^{\perp}$ are those in $\mathrm{Im}(f)$. This means that since $\mathcal{C}=\mathrm{Im}(f)^{\perp}/\mathrm{Im}(f)$ we must have that braiding is nondegenerate on $\mathcal{C}$.